# definitionDefinition

### v1

@@ I Introduction
The rest of this paper is organized as follows: Sec. II gives an overview of relevant background material on property types, DNN verifiers, and NLB certificates. Related work is covered in Sec. III. In Sec. IV and V, we present our approach, and Sec. VI reports the results of our spacecraft case study. Finally, Sec. VII concludes.
@@ III-A1 Formal Verification of Neural Certificates
Recent methods have leveraged neural networks as verifiable models of these control certificates, forming a class of neural certificate approaches [30]. For a fixed controller, [68] distills the problem into solving binary classification with neural networks, but the method is limited to polynomial systems and only obtains a region of attraction.
@@ IV-A RWA certificates
If V is an RWA certificate for a dynamical system with witness $(\alpha,\beta,\epsilon)$, then for every trajectory $\tau$ starting from a state $x\in\mathcal{X}\setminus\mathcal{X}_{G}$ such that $V(x)\leq\beta$, $\tau$ will eventually contain a state in $\mathcal{X}_{G}$ without ever passing through a state in $\mathcal{X}_{U}$.
@@ IV-B FRWA certificates
Eq. (7) penalizes deviations from constraint (4), and Eq. (8) penalizes deviations from constraint (5). We incorporate parameters $\delta_{1}>0$ and $\delta_{2}>0$, which can be used to tune how strongly the certificate over-approximates adherence to each constraint. Similarly, constants $c_{s}$ and $c_{d}$ can be used to tune the relative weight of the two objectives. The final training objective $O$ in (9) is what the optimizer seeks to minimize, by using stochastic gradient descent (SGD) or other optimization techniques.
@@ IV-C CEGIS loop
We use a counterexample-guided inductive synthesis (CEGIS) loop, shown in Fig. 1, to obtain a fully verified controller and certificate. We first train an initial controler $\pi$. Then, at each CEGIS iteration, we jointly train $V$ and $\pi$ until a loss of 00 is obtained and then use a sound and complete DNN verifier (we use Marabou [51] in our experiments) to identify counterexamples. If the verifier identifies a counterexample violating constraints (4) or (5) (recall that constraint (6) is satisfied by construction), we sample points in the proximity of the counterexample and use these to augment the training data. By sampling multiple nearby points, we hope to influence the training to learn smooth behavior for a localized neighborhood instead of overfitting to a specific point. This process is repeated iteratively until no counterexamples are found, at which point we are guaranteed to have produced a fully verified controller and certificate.
@@ V Compositional Certificates
Formally, a compositional RWA certificate (CRWA) for an RWA task is composed of $n$ RWA certificates, which we denote $V_{0},\dots,V_{n-1}$, with corresponding controllers, which we denote $\pi_{0},\dots,\pi_{n-1}$, with $n\geq 2$. Furthermore, each pair $(V_{i},\pi_{i})$ must be an RWA certificate with some witness $(\alpha_{i},\beta_{i},\epsilon_{i})$ for an RWA task whose dynamics are that of the main RWA task, but whose parameters are $(\mathcal{X}^{i}_{I},\mathcal{X}^{i}_{G},\mathcal{X}^{i}_{U})$. These parameters must satisfy the following conditions:
@@
    $\mathcal{X}^{n}_{I}=\mathcal{X}_{I}$ and $\mathcal{X}^{n}_{U}=\mathcal{X}_{U}$.
Intuitively, the idea is as follows. We start with an initial controller capable of guiding the system from some subset of the initial states $\mathcal{X}_{I}$ to the goal states $\mathcal{X}_{G}$ while avoiding some superset of the unsafe states $\mathcal{X}_{U}$. Then, for each subsequent controller, we ensure that it can guide the system either from a larger subset of the initial states $\mathcal{X}_{I}$ or while avoiding a smaller superset of the unsafe states $\mathcal{X}_{U}$, or both, to a new goal region consisting of the states considered safe by the previous controller, i.e., the states $x$ for which $V(x)\leq\beta$. For the final controller (controller $n-1$), the set of initial and unsafe states should coincide with those of the original RWA problem.
@@ 10       Choose $\alpha^{i}>\beta^{i}$ and $\epsilon^{i}>0$
Figure 2: Visualization of how consecutive RWA tasks relate when building a CRWA certificate.
The training and verification of a CRWA certificate is described in Alg. 1 and visualized in Fig. 2. The following lemma captures the correctness of our approach. [^3] Due to space constraints, the proof of this lemma is in Appendix III.
@@
When training certificate $V_{i}$, it is important that the training dataset contains sufficient states sampled from $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$. Otherwise, $V_{i}$ might learn to assign values greater than $\beta^{i}$ as much as possible in order to meet constraint (5). However, this cannot happen for states in $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$ because of constraint (4). To ensure that states in the region $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$ are included in the training data, we can query a DNN verifier to identify points in that region and then include in the data set those points as well as a random subset of their neighbors.
Initialization for CRWA Training.
When training certificate $V_{i}$, it is important that the training dataset contains sufficient states sampled from $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$. Otherwise, $V_{i}$ might learn to assign values greater than $\beta^{i}$ as much as possible in order to meet constraint (5). However, this cannot happen for states in $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$ because of constraint (4). To ensure that states in the region $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$ are included in the training data, we can query a DNN verifier to identify points in that region and then include in the data set those points as well as a random subset of their neighbors.
@@
$\mathcal{X}^{i}_{G}=\{x\in\overline{\mathcal{X}^{i-1}_{U}}\ \mid\ V_{i-1}(x)<\beta_{i-1}\}\cup\mathcal{X}^{i-1}_{G}.$ (11)
@@ VI-A Case Study
We evaluate our approach on the 2D docking task from [66], in which a spacecraft is trained using DRL to navigate to a goal. More specifically, a DRL agent maneuvers a deputy spacecraft, controlled with thrusters that provide forces in the $x$ and $y$ directions. The deputy spacecraft attempts to safely navigate until it reaches a state that is in close proximity to a designated chief spacecraft, while obeying a distance-dependent safety constraint. [^4] Further details are in Appendix II. We focus on this benchmark for several reasons: (i) it has been proposed and studied as a challenge problem in the literature [66], (ii) there exist natural safety and liveness properties for it; and (iii) existing approaches have been been unable to formally verify these properties.
@@
The goal is for the deputy to successfully dock with the chief without ever violating the velocity safety constraint. In the original benchmark [66], the docking (goal) region is defined as a circle of radius $r=0.5$m around the origin $(0,0)$. In our evaluation, we use this same goal region for the initial training of the DRL controller. However, during the CEGIS iteration (i.e., the “Train” and “Verify” steps in Fig. 1), we use a conservative subset of this goal region, namely a square centered at the origin whose sides have length $l=0.7$m. The reason for this is so that the goal region can easily be described using linear inequalities.
@@ VI-C Implementation and Setup
All of our experiments aim to solve RWA tasks, as defined in Definition II.1. The system dynamics are those of the 2D spacecraft, as described in Section VI-A. RWA tasks are parameterized by $\mathcal{X}_{G}$, $\mathcal{X}_{I}$, and $\mathcal{X}_{U}$. For these sets of states, we typically use square regions centered at the origin. For convenience, we refer to the set $\{(x,y)\ \mid\ x,y\in[-a,a]\}$ with the abbreviation $[-a,a]$. For example, as outlined in Section VI-A, we set $\mathcal{X}_{G}=[-0.35,0.35]$. We use different values for $\mathcal{X}_{I}$, depending on the experiment (indeed, this is the primary variable we vary in our experiments), but whenever $\mathcal{X}_{I}=[-a,a]$, we then set $\mathcal{X}_{U}=[-(a+1),a+1]$.
@@ VI-D Experimental Results
TABLE II: Compositional certificate results. The columns indicate: the initialization set for the final certificate, the size of the compositional certificate, the initialization set for all but the final certificate, the cumulative total time for all but the final certificate, and the statistics for training the final certificate.
@@
Each row of the table corresponds to a compositional certificate. The first column shows the value of $\mathcal{X}_{I}$ for this certificate. The next columns indicate the size $n$ of the certificate, the values of $\mathcal{X}^{i}_{I}$ for $10\leq i<n-$ 1, and the cumulative time required for all but the last certificate. The next three columns give the minimum, mean, and maximum time required to produce the controller and certificate for the last stage of the compositional certificate (recall that we run five independent trials for all CEGIS loops). The next three columns show the minimum, mean, and maximum number of CEGIS iterations used, and the last column indicates how many of the trials succeeded.
The results clearly indicate that compositional certificates greatly improve scalability. Whereas the stand-alone certificates could not scale beyond $[-3,3]$ in 12 hours, we were able to successfully produce a formally verified $5$-stage certificate for $[-11,11]$ in a little over 9 hours. It is also worth noting that we do get a significant benefit by running 5 independent CEGIS loops, as both the time and the number of loops can vary significantly from the minimum to the maximum. Nearly all of the CEGIS loops eventually completed—only the initial $[3,3]$ region failed to complete all of its trials—suggesting that the compositional approach is also more stable and robust. This can also be seen in Fig. 4: for each starting region $[-a,a]$, the diamond point labeled CFRWA-$i$ corresponds to the $i$th row containing $[-a,a]$ in column 1. We can see that, compared to the stand-alone RWA and FRWA certificates, the compositional certificates can be trained faster and with fewer failures.
@@ VII Conclusion
In the future, we plan to extend our approach to be compatible with additional formal techniques (e.g., shielding). We also plan to apply our approach to more challenging case studies with larger DRL controllers. We see this work as an important step towards the safe and reliable use of DRL in real-world systems.
@@ II-A3 Observations
: The observations of both (i) the scalar speed, and (ii) the distance dependent velocity limit, were removed. This is due to their non-linear characterization with regard to the system’s state space. However, the x, y position and velocity features were retained in the agent’s observation space.
@@ III Proofs
If V is an RWA certificate for a dynamical system with witness $(\alpha,\beta,\epsilon)$, then for every trajectory $\tau$ starting from a state $x\in\mathcal{X}\setminus\mathcal{X}_{G}$ such that $V(x)\leq\beta$, $\tau$ will eventually contain a state in $\mathcal{X}_{G}$ without ever passing through a state in $\mathcal{X}_{U}$.

### v2

@@ I Introduction
The rest of this paper is organized as follows: Sec. II gives an overview of relevant background material on property types, DNN verifiers, and NLB certificates. Related work is covered in Sec. III. In Sec. IV and V, we present our approach, and Sec. VI reports the results of our spacecraft case study. [^1] Code for reproducing experiments is available at: github.com/NeuralNetworkVerification/artifact-fmcad24-docking Finally, Sec. VII concludes.
@@ III-A1 Formal Verification of Neural Certificates
Recent methods have leveraged neural networks as verifiable models of these control certificates, forming a class of neural certificate approaches [34]. For a fixed controller, [80] distills the problem into solving binary classification with neural networks, but the method is limited to polynomial systems and only obtains a region of attraction, making it incompatible with most RWA problems, which have a predefined goal region.
@@ IV-A RWA certificates
If V is an RWA certificate for a dynamical system with witness $(\alpha,\beta,\epsilon)$, and $V$ has a lower bound, [^4] This is always the case if the output of $V$ is implemented using a finite representation such as floating-point arithmetic. then for every infinite trajectory $\tau$ starting from a state $x\in\mathcal{X}\setminus\mathcal{X}_{G}$ such that $V(x)\leq\beta$, $\tau$ will eventually contain a state in $\mathcal{X}_{G}$ without ever passing through a state in $\mathcal{X}_{U}$.
@@ IV-B FRWA certificates
Eq. (7) penalizes deviations from constraint (4), and Eq. (8) penalizes deviations from constraint (5). We incorporate parameters $\delta_{1}>0$ and $\delta_{2}>0$, which can be used to tune how strongly the certificate over-approximates adherence to each constraint. Similarly, constants $c_{s}$ and $c_{d}$ can be used to tune the relative weight of the two objectives. The final training objective $O$ in (9) is what the optimizer seeks to minimize, by using stochastic gradient descent (SGD) or other optimization techniques. We note that the FRWA certificates are trained in a self-supervised, non-RL setting.
@@ IV-C CEGIS loop
We use a counterexample-guided inductive synthesis (CEGIS) loop, shown in Fig. 1, to obtain a fully verified controller and certificate. We first train an initial controller $\pi$. Then, at each CEGIS iteration, we jointly train $V$ and $\pi$ until a loss of 00 is obtained and then use a sound and complete DNN verifier (we use Marabou [59] in our experiments) to identify counterexamples. If the verifier identifies a counterexample violating constraints (4) or (5) (recall that constraint (6) is satisfied by construction), we sample points in the proximity of the counterexample and use these to augment the training data. By sampling multiple nearby points, we hope to influence the training to learn smooth behavior for a localized neighborhood instead of overfitting to a specific point. This process is repeated iteratively until no counterexamples are found, at which point we are guaranteed to have produced a fully verified controller and certificate.
@@ V Compositional Certificates
Formally, a compositional RWA certificate (CRWA) for an RWA task is composed of $n$ RWA certificates, [^5] Each controller in a CRWA can make use of the FRWA technique described above. which we denote $V_{0},\dots,V_{n-1}$, with corresponding controllers, which we denote $\pi_{0},\dots,\pi_{n-1}$, with $n\geq 2$. Furthermore, each pair $(V_{i},\pi_{i})$ must be an RWA certificate with some witness $(\alpha_{i},\beta_{i},\epsilon_{i})$ for an RWA task whose dynamics are that of the main RWA task, but whose parameters are $(\mathcal{X}^{i}_{I},\mathcal{X}^{i}_{G},\mathcal{X}^{i}_{U})$. These parameters must satisfy the following conditions:
@@
    $\mathcal{X}^{n-1}_{I}=\mathcal{X}_{I}$ and $\mathcal{X}^{n-1}_{U}=\mathcal{X}_{U}$.
Intuitively, the idea is as follows. We start with an initial controller capable of guiding the system from some subset of the initial states $\mathcal{X}_{I}$ to the original goal states $\mathcal{X}_{G}$ while avoiding some superset of the unsafe states $\mathcal{X}_{U}$. Then, for each subsequent controller, we ensure that it can guide the system either from a larger subset of the initial states $\mathcal{X}_{I}$ or while avoiding a smaller superset of the unsafe states $\mathcal{X}_{U}$, or both, to a new goal region consisting of the states considered safe by the previous controller, i.e., the states $x$ for which $V(x)\leq\beta$. For the final controller (controller $n-1$), the set of initial and unsafe states should coincide with those of the original RWA problem. Note that the algorithm does not say how to choose of $\mathcal{X}_{I}^{i}$ and $\mathcal{X}_{U}^{i}$ for $i<n-1$ other than to specify that these sets should be monotonically increasing and decreasing, respectively. Finding good heuristics for choosing these sets in the general case is a promising direction for future work.
@@ 10       Choose $\alpha^{i}>\beta^{i}$ and $\epsilon^{i}>0$
Figure 2: Visualization of how consecutive certificates relate when building a CRWA certificate. Note that $\mathcal{X}_{G}$ need not be a subset of $\mathcal{X}_{I}$. The dotted lines indicate that the unsafe state region extends infinitely outside the solid line box. Wavy lines indicate outer boundaries for initial or goal regions.
The training and verification of a CRWA certificate is described in Alg. 1 and visualized in Fig. 2.
The following lemma captures the correctness of our approach. [^6] Due to space constraints, the proof of this lemma is in Appendix III.
@@
When training certificate $V_{i}$, it is important that the training dataset contains sufficient states sampled from $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$. Otherwise, $V_{i}$ might learn to assign values greater than $\beta^{i}$ as much as possible in order to meet constraint (5), as opposed to appropriately assigning all states in $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$ to have values less than $\beta^{i}$, due to an insufficient loss penalty for constraint (6). To ensure that states in the region $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$ are included in the training data, we can identify states over constrained subspaces in $\mathcal{X}^{i}_{I}\setminus\mathcal{X}^{i}_{G}$, and then include in the data set those points as well as a random subset of their neighbors which likely lie in the same region.
@@
$\mathcal{X}^{i}_{G}=\{x\in\overline{\mathcal{X}^{i-1}_{U}}\ \mid\ V_{i-1}(x)\leq\beta_{i-1}\}\cup\mathcal{X}^{i-1}_{G}.$ (11)
@@ VI-A Case Study
We evaluate our approach on the 2D docking task from [78], [^7] Additional details on this case study are described in our recent related paper [70] in which a spacecraft is trained using DRL to navigate to a goal. More specifically, a DRL agent maneuvers a deputy spacecraft, controlled with thrusters that provide forces in the $x$ and $y$ directions. The deputy spacecraft attempts to safely navigate until it reaches a state that is in close proximity to a designated chief spacecraft, while obeying a distance-dependent safety constraint. [^8] Further details are in Appendix II. We focus on this benchmark for several reasons: (i) it has been proposed and studied as a challenge problem in the literature [78], (ii) there exist natural safety and liveness properties for it; and (iii) existing approaches have been been unable to formally verify these properties.
@@
The goal is for the deputy to successfully dock with the chief without ever violating the velocity safety constraint. In the original benchmark [78], the docking (goal) region is defined as a circle of diameter $d=1$m centered at the origin $(0,0)$. In our evaluation, we use this same goal region for the initial training of the DRL controller. However, during the CEGIS iteration (i.e., the “Train” and “Verify” steps in Fig. 1), we use a conservative subset of this goal region, namely a square centered at the origin whose sides have length $l=0.7$m. The reason for this is so that the goal region can easily be described using linear inequalities.
@@ VI-C Implementation and Setup
All of our experiments aim to solve RWA tasks, as defined in Definition II.1. The system dynamics are those of the 2D spacecraft, as described in Section VI-A. RWA tasks are parameterized by $\mathcal{X}_{G}$, $\mathcal{X}_{I}$, and $\mathcal{X}_{U}$. For these sets of states, we typically use square regions centered at the origin. For convenience, we refer to the set $\{(x,y)\ \mid\ x,y\in[-a,a]\}$ with the abbreviation $[-a,a]$. For example, as outlined in Section VI-A, we set $\mathcal{X}_{G}=[-0.35,0.35]$. We use different values for $\mathcal{X}_{I}$, depending on the experiment (indeed, this is the primary variable we vary in our experiments), but whenever $\mathcal{X}_{I}=[-a,a]$, we then set $\overline{\mathcal{X}_{U}}=[-(a+1),a+1]$.
@@ VI-D Experimental Results
TABLE II: Compositional certificate results. The columns indicate: the initial set for the final certificate, the size of the compositional certificate, the initial sets for all but the final certificate, the cumulative time for all but the final certificate, the total time (min, mean, and max) for training all certificates, and statistics for training the final certificate. We note that the cumulative time column is always equal to the corresponding value in the min column corresponding to the penultimate certificate. The wall time is the total time including the final controller/certificate. The CEGIS iterations and success stats are for the final controller/certificate only.
@@
Each row of the table corresponds to a compositional certificate. The first column shows the value of $\mathcal{X}_{I}$ for this certificate. The next columns indicate the number $n$ of composed certificates, the values of $\mathcal{X}^{i}_{I}$ for $10\leq i<n-$ 1, and the cumulative time required for all but the last certificate. The next three columns give the minimum, mean, and maximum time required to produce the controller and certificate for the last stage of the compositional certificate (recall that we run five independent trials for all CEGIS loops). The next three columns show the minimum, mean, and maximum number of CEGIS iterations used, and the last column indicates how many of the trials succeeded. Note that when $n=1$, the row corresponds to a single FRWA certificate.
The results clearly indicate that compositional certificates greatly improve scalability. Whereas the stand-alone certificates could not scale beyond $[-3,3]$ in 12 hours, we were able to successfully produce a formally verified $5$-stage certificate for $[-11,11]$ in a little over $5.7$ hours. It is also worth noting that we do get a significant benefit by running 5 independent CEGIS loops, as both the time and the number of loops can vary significantly from the minimum to the maximum. Nearly all of the CEGIS loops eventually completed—only the initial $[3,3]$ region failed to complete all of its trials—suggesting that the compositional approach is also more stable and robust. This can also be seen in Fig. 4: for each starting region $[-a,a]$, the diamond point labeled CFRWA-$i$ corresponds to the $i$th row containing $[-a,a]$ in column 1. We can see that, compared to the stand-alone RWA and FRWA certificates, the compositional certificates can be trained faster and with fewer failures.
@@ VII Conclusion
In the future, we plan to extend our approach to be compatible with additional formal techniques (e.g., shielding against safety violations [89, 18, 6, 74, 61, 32, 81], and Scenario-Based Programming [45, 58, 48, 60, 55, 33, 92, 93]). We also plan to apply our approach to more challenging case studies with larger DRL controllers. We see this work as an important step towards the safe and reliable use of DRL in real-world systems.
@@ II-A3 Observations
: The observations of both (i) the scalar speed, and (ii) the distance-dependent velocity limit, were removed. This is due to their non-linear characterization with regard to the system’s state space. However, the x, y position and velocity features were retained in the agent’s observation space. It is important to note that the distance-dependent velocity limit can be reintroduced and learned through a post-training process, like the RWA training defined in our work.
@@ III Proofs
If V is an RWA certificate for a dynamical system with witness $(\alpha,\beta,\epsilon)$, and $V$ has a lower bound, then for every infinite trajectory $\tau$ starting from a state $x\in\mathcal{X}\setminus\mathcal{X}_{G}$ such that $V(x)\leq\beta$, $\tau$ will eventually contain a state in $\mathcal{X}_{G}$ without ever passing through a state in $\mathcal{X}_{U}$.